Circular No. 7202

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cn IN MCG +2-38-043 S. Benetti, Telescopio Nazionale Galileo (TNG); C. Morossi, Osserva-

torio Astronomico di Trieste; and F. Bortoletto, R. Cosentino, D. Gardiol, A. Ghedina, F. Ghinassi, A. Magazzú, E. Marchetti, C. Pernechele, and A. Zacchei, TNG, report the discovery of a supernova (V ∼ 17.9) on a V CCD frame taken with the TNG (+ Optical Imager Galileo) at La Palma on June 14.99 UT, while observing the galaxy MCG +2-38-043 (CGCG 077-007; α = 15^{h}06^{m}46^{s}.6, δ = +12^{o}51′29′′, equinox 2000.0). SN 1999cn is located on a spiral arm and its offset from the galaxy center is 1′′.5 west and 8′′.3 north. Benetti also reports that an inspection of a preliminarily reduced spectrum (range 328–760 nm; resolution 0.35 nm) taken at the William Herschel Telescope (+ ISIS) by D. Axon, University of Hertfordshire; C. Packham, A. Humphrey, and J. Ray, Isaac Newton Group; and J. Smith, University of Hertfordshire, on June 15.14 shows that the supernova is of type Ic near maximum. The spectrum is dominated by strong P-Cyg lines of Ca II (H and K) and Fe II. A weaker Si II 635.5-nm line is also visible. The expansion velocities as derived from Ca II and Si II lines are ∼ 13 900 and ∼ 11 300 km/s, respectively. The redshift of the galaxy, derived from a strong and narrow Hα emission superimposed on the supernova spectrum, is ∼ 6790 km/s.

XTE J1550–564 Correcting their report on IAUC 7184, R. Soria and K. Wu write: “The

G-star spectrum that was provided by the Anglo-Australian Observatory through a service observation on Apr. 10 is not the spectrum of the secondary star of RXTE J1550–564. Therefore, the spectral type and distance estimate based on this spectroscopic observation are invalid.”

V854 CENTAURI This R CrB-type variable has undergone a rapid fading, as indicated by

the following visual magnitude estimates: June 4.56 UT, 7.8 (P. Williams, Heathcote, N.S.W.); 6.35, 8.0 (Williams); 6.98, 8.0 (A. Pereira, Cabo da Roca, Portugal); 7.38, 8.6 (Williams); 7.94, 8.3 (Pereira); 11.47, 10.9 (Williams); 13.44, 11.9 (Williams).

1999 June 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green